OpenStudy (anonymous):

Joseph borrowed a book from a library. The library charged a fixed rental for the book and a late fee for every day the book was overdue. The expression below shows the charges Joseph paid for the book when he returned it x days after the due date. 2 + 0.25x What does the coefficient of the expression represent? The fixed rental for the book The total late fee for the book The late fee per day for the book The number of days the book was overdue

OpenStudy (anonymous):

i think its c bby

OpenStudy (anonymous):

thanks! An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

im pretty sure

OpenStudy (anonymous):

Evaluate: -6 - (-3) 18 -9 3 -3

OpenStudy (anonymous):

d